# Video Compression Commander: Plug-and-Play Inference Acceleration for Video Large Language Models

### v1

@@ Appendix A Benchmark Details
  • MVBench (Li et al., 2024b) defines 20 video understanding tasks that require deep comprehension of temporal dimensions, beyond single-frame analysis.
  • LongVideoBench (Wu et al., 2024) focuses on long-context video understanding with 3,763 videos up to one hour long. It includes 6,678 multiple-choice questions across 17 categories, emphasizing temporal information retrieval and analysis.
  • MLVU (Zhou et al., 2024) features videos ranging from 3 minutes to 2 hours, encompassing 9 evaluation tasks including topic reasoning, anomaly recognition, video summarization, and plot question-answering.
  • VideoMME (Fu et al., 2024) comprises 900 videos and 2,700 multiple-choice questions across six domains, with durations from 11 seconds to 1 hour, categorized into short, medium, and long subsets.
@@ Appendix B Model Details
  • LLaVA-OneVision (Li et al., 2024a) unifies single-image, multi-image, and video tasks in a single LLaVA-OneVision model. It represents videos as long visual token sequences in the same “interleaved” format used for images, enabling smooth task transfer from images to videos and facilitating strong zero-shot video understanding capabilities.
  • LLaVA-Video (Zhang et al., 2024c) builds upon the single-image stage checkpoint of LLaVA-OneVision. It is fine-tuned on a large synthetic video-instruction dataset (LLaVA-Video-178K), covering detailed captioning, open-ended QA, and multiple-choice QA. By employing the SigLIP visual encoder and Qwen2 as the LLM, LLaVA-Video achieves robust video comprehension across various benchmarks.
  • Qwen2-VL (Wang et al., 2024) introduces Naive Dynamic Resolution to adaptively convert frames of any resolution into visual tokens. It utilizes Multimodal Rotary Position Embedding within a unified image-and-video processing paradigm, enabling the handling of long videos (20+ minutes) for high-quality QA, dialogue, and content creation.
@@ Appendix C Baseline Details
  • FastV (Chen et al., 2024a) performs one-time token pruning as an intra-LLM compression method, utilizing attention weights associated with the output token after a selected LLM layer. However, its explicit dependence on attention weights makes it incompatible with Flash Attention (Dao et al., 2022) in LLM.
  • PDrop (Xing et al., 2025) extends intra-LLM compression by implementing progressive token pruning across multiple LLM layers, based on attention weights of output tokens. Similarly, this explicit attention mechanism prevents compatibility with Flash Attention (Dao et al., 2022) in LLM.
  • SparseVLM (Zhang et al., 2025) functions as an intra-LLM compression method, ranking token importance using text-visual attention maps and pruning via pre-selected text prompts to mitigate attention noise. Similar to FastV, SparseVLM is also incompatible with Flash Attention (Dao et al., 2022) in LLM.
  • MUSTDrop (Liu et al., 2024) is a three-stage compression method operating in ViT and LLM stages. It relies on [CLS] token attention and text-visual attention for token selection and pruning. This approach faces compatibility issues with [CLS]-free VideoLLMs and prevents Flash Attention support in LLM due to its explicit use of attention weights.
  • FiCoCo (Han et al., 2024) is a two-stage compression method that merges tokens in ViT using [CLS] and patch-patch attention, then further compresses in LLM using text-visual attention. It suffers from [CLS] dependency and lacks Flash Attention compatibility.
  • FasterVLM (Zhang et al., 2024b) is another pre-LLM compression method that relies on [CLS] token attention weights to retain informative visual tokens. It also faces compatibility issues with [CLS]-free VideoLLMs and Flash Attention integration in ViT.
  • DyCoke (Tao et al., 2025) is a two-stage VideoLLM-specific method that first prunes similar tokens along the temporal dimension and then uses attention weights in the LLM to compress the less attended visual tokens in the KV cache. Due to its reliance on dividing frame sets into parts and compressing them through similarity calculations, similar to ToMe (Bolya et al., 2023), it cannot achieve aggressive token compression in one go. While its token compression stage is compatible with Flash Attention (Dao et al., 2022), its KV cache compression requires explicit attention weights and thus remains incompatible with efficient attention operators.

### v2

@@ 6 Limitations
Acknowledgement
This research was supported by the Shanghai Science and Technology Program (Grant No. 25ZR1402278). Besides, we thank Huawei Ascend Cloud Ecological Development Project for the support of Ascend 910 processors.
@@ Appendix A Benchmark Details
  • MVBench Li et al. (2024b) defines 20 video understanding tasks that require deep comprehension of temporal dimensions, beyond single-frame analysis.
  • LongVideoBench Wu et al. (2024) focuses on long-context video understanding with 3,763 videos up to one hour long. It includes 6,678 multiple-choice questions across 17 categories, emphasizing temporal information retrieval and analysis.
  • MLVU Zhou et al. (2024) features videos ranging from 3 minutes to 2 hours, encompassing 9 evaluation tasks including topic reasoning, anomaly recognition, video summarization, and plot question-answering.
  • VideoMME Fu et al. (2024) comprises 900 videos and 2,700 multiple-choice questions across six domains, with durations from 11 seconds to 1 hour, categorized into short, medium, and long subsets.
  • EgoSchema Mangalam et al. (2023) consists of 5,000 egocentric videos with multiple-choice questions requiring comprehensive understanding of procedural activities and temporal reasoning over extended sequences, challenging models with first-person perspective video analysis.
  • PerceptionTest Patraucean et al. (2023) presents 11,609 real-world videos with 38,565 multiple-choice questions evaluating diverse perceptual skills including object tracking, action recognition, and temporal localization across varied scenarios and contexts.
@@ Appendix B Model Details
  • LLaVA-OneVision Li et al. (2024a) unifies single-image, multi-image, and video tasks in a single LLaVA-OneVision model. It represents videos as long visual token sequences in the same “interleaved” format used for images, enabling smooth task transfer from images to videos and facilitating strong zero-shot video understanding capabilities.
  • LLaVA-Video Zhang et al. (2024c) builds upon the single-image stage checkpoint of LLaVA-OneVision. It is fine-tuned on a large synthetic video-instruction dataset (LLaVA-Video-178K), covering detailed captioning, open-ended QA, and multiple-choice QA. By employing the SigLIP visual encoder and Qwen2 as the LLM, LLaVA-Video achieves robust video comprehension across various benchmarks.
  • Qwen2-VL Wang et al. (2024) introduces Naive Dynamic Resolution to adaptively convert frames of any resolution into visual tokens. It utilizes Multimodal Rotary Position Embedding within a unified image-and-video processing paradigm, enabling the handling of long videos (20+ minutes) for high-quality QA, dialogue, and content creation.
@@ Appendix C Baseline Details
  • FastV Chen et al. (2024a) performs one-time token pruning as an intra-LLM compression method, utilizing attention weights associated with the output token after a selected LLM layer. However, its explicit dependence on attention weights makes it incompatible with Flash Attention Dao et al. (2022) in LLM.
  • PDrop Xing et al. (2025) extends intra-LLM compression by implementing progressive token pruning across multiple LLM layers, based on attention weights of output tokens. Similarly, this explicit attention mechanism prevents compatibility with Flash Attention Dao et al. (2022) in LLM.
  • SparseVLM Zhang et al. (2025) functions as an intra-LLM compression method, ranking token importance using text-visual attention maps and pruning via pre-selected text prompts to mitigate attention noise. Similar to FastV, SparseVLM is also incompatible with Flash Attention Dao et al. (2022) in LLM.
  • MUSTDrop Liu et al. (2024) is a three-stage compression method operating in ViT and LLM stages. It relies on [CLS] token attention and text-visual attention for token selection and pruning. This approach faces compatibility issues with [CLS]-free VideoLLMs and prevents Flash Attention support in LLM due to its explicit use of attention weights.
  • FiCoCo Han et al. (2024) is a two-stage compression method that merges tokens in ViT using [CLS] and patch-patch attention, then further compresses in LLM using text-visual attention. It suffers from [CLS] dependency and lacks Flash Attention compatibility.
  • FasterVLM Zhang et al. (2024b) is another pre-LLM compression method that relies on [CLS] token attention weights to retain informative visual tokens. It also faces compatibility issues with [CLS]-free VideoLLMs and Flash Attention integration in ViT.
  • DyCoke Tao et al. (2025) is a two-stage VideoLLM-specific method that first prunes similar tokens along the temporal dimension and then uses attention weights in the LLM to compress the less attended visual tokens in the KV cache. Due to its reliance on dividing frame sets into parts and compressing them through similarity calculations, similar to ToMe Bolya et al. (2023), it cannot achieve aggressive token compression in one go. While its token compression stage is compatible with Flash Attention Dao et al. (2022), its KV cache compression requires explicit attention weights and thus remains incompatible with efficient attention operators.